A body of mass 40g runs at a speed of 6km/h what is it kinectic energy?



1
Expert's answer
2021-01-15T13:46:41-0500

Be the definition of the kinetic energy, we have:


"KE=\dfrac{1}{2}mv^2,""KE=\dfrac{1}{2}\cdot0.04\ kg\cdot(6\ \dfrac{km}{h}\cdot \dfrac{1000\ m}{1\ km}\cdot \dfrac{1\ h}{3600\ s})^2=0.055\ J."

Answer:

"KE=0.055\ J."